A distribution centre sends three different types of parcels. One consignment has 6 small, 8
medium and 9 large parcels which cost $173.20 to post. Another consignment has 7 small, 13
medium and 17 large parcels with total postage $291.05. A large parcel costs twice as much to
post as a small one. Work out the total cost of posting 3 small parcels, 9 medium parcels and 2
large parcels.

Answer by greenestamps(13200)



x = cost for small parcel
y = cost for medium parcel
2x = cost for large parcel

6 small, 8 medium, 9 large:
6(x)+8(y)+9(2x) = 173.20
24x+8y = 173.20 (1)

7 small, 13 medium, 17 large:
7(x)+13(y)+17(2x) = 291.05
41x+13y = 291.05 (2)

Simplify (1) to 3x+y = 21.65 (3)

Multiply by 13: 39x+13y = 281.45 (4)

Compare (2) and (4): 2x = 9.60 --> x = 4.80

Plug x = 4.80 into any earlier equation to find y = 7.25

ANSWER:
3(x)+9(y)+2(2x) = 7x+9y = 7(4.80)+9(7.25) = 98.85